# Decentralized real-time iterations for distributed NMPC

### v2

@@ III Centralized Real-Time Iterations
Furthermore, there exists a constant $L_{V,x}>0$ such that $|\sqrt{V(x)}-\sqrt{V(x^{\prime})}|\leq L_{V,x}\|x-x^{\prime}\|$ for all $x,x^{\prime}\in\mathbb{X}_{\bar{V}}.\hfill\square$
@@
Let $\bar{p}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{p}}$ be the map from the current state $x(t)$ in (2c) to the globally optimal primal-dual variables $p^{\star}$ of OCP (2). $\square$
@@
Assumption 2 (Lipschitz controller [Zanelli2021])
There exist positive constants $\hat{r}_{x}$ and $L_{p,x}$ such that $\|\bar{p}(x^{\prime})-\bar{p}(x)\|\leq L_{p,x}\|x^{\prime}-x\|$ for all $x\in\mathbb{X}_{\bar{V}}$ and for all $x^{\prime}\in\mathcal{B}(x,\hat{r}_{x})$. Moreover, $\bar{p}(0)=0$. $\square$
Note that $\bar{p}(0)=0$ if the origin is an equilibrium of the system dynamics which minimizes $\ell$ and $V_{\mathrm{f}}$, a common design choice in stabilizing NMPC. Put differently, this assumption does not hold for so-called economic NMPC schemes without further modification. The superscript $\cdot^{k}$ denotes the iteration index of the optimization method which is used to solve OCP (2).
Assumption 3 (Q-linear optimizer convergence [Zanelli2021])
There exist positive constants $\hat{r}_{p}>0$ and $a_{p}<1$ such that, for all $x\in\mathbb{X}_{\bar{V}}$ and for all $p^{k}\in\mathcal{B}(\bar{p}(x),\hat{r}_{p})$, the optimization method initialized with $p^{k}$ returns an iterate $p^{k+1}$ satisfying
$\left\|p^{k+1}-\bar{p}(x)\right\|\leq a_{p}\left\|p^{k}-\bar{p}(x)\right\|.$
@@
Assumption 4 (Lipschitz system dynamics [Zanelli2021])
The centralized dynamics satisfy $f^{\mathrm{c}}(0,0)=0$. Consider $r_{p}$ from Lemma 7 in Appendix A. There exist positive finite constants $\varrho$, $L_{f,x}^{\mathrm{c}}$, and $L_{f,u}^{\mathrm{c}}$ such that
@@
for all $x^{\prime},x\in\mathbb{X}_{\bar{V}}\oplus\mathcal{B}(0,\varrho)$ and for all $u^{\prime}=M_{u,p}p^{\prime}$, $u=M_{u,p}p$ with $p^{\prime},p\in\mathcal{B}(\bar{p}(x),{r}_{p})$. $\square$
At time $t$, the state $x(t)$ is sampled and the optimization method is initialized with the solution from the previous control step, i.e., $p^{0}(t)=p^{k_{\text{max}}}(t-1)$, where $k_{\text{max}}\in\mathbb{N}$ is the number of optimizer iterations per control step. Then, $k_{\text{max}}$ optimizer iterations are applied to the OCP, a primal-dual iterate $p^{k_{\text{max}}}(t)$ is obtained, and the control input is selected. The control input is applied to the system and the primal-dual iterates are stored for the next time step. The optimization method and the system together form the system-optimizer dynamics [Zanelli2021]
$\begin{bmatrix}x(t+1)\\ p^{k_{\text{max}}}(t+1)\end{bmatrix}=\begin{bmatrix}f^{\delta}(x(t),M_{u,p}p^{k_{\text{max}}}(t))\\ \Phi(x(t),p^{k_{\text{max}}}(t))\end{bmatrix},$ (4)
where $\Phi:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{p}}\rightarrow\mathbb{R}^{n_{p}}$ maps the OCP initial state $x(t)$ and approximate solution $p^{k_{\text{max}}}(t)$ to the approximate solution $p^{k_{\text{max}}}(t+1)$ at the next time step.
Lemma 1 (Centralized RTI stability [Zanelli2021])
Suppose that As-sumptions 14 hold and consider the sufficient sampling interval $\bar{\delta}$ and optimizer initialization radius $\tilde{r}_{p}$ defined in (29) in Appendix A. If $\delta\leq\bar{\delta}$, $x(0)\in\mathbb{X}_{\bar{V}}$, and $p^{0}(0)\in\mathcal{B}(\bar{p}(x(0)),\tilde{r}_{p})$, then the origin is an exponentially stable equilibrium for the closed-loop system-optimizer dynamics (4). $\square$
Remark 1 (Relation to [Zanelli2021])
Lemma 1 is a specialized variant of [Zanelli2021, Thm. 25] in three aspects: First, we only consider quadratic bounds on the Lyapunov function. Second, we define the map $\bar{p}$ from the current state to the primal-dual variables, because of the dSQP convergence presented in the next section. Third, we allow for $k_{\mathrm{max}}\geq 1$ optimizer iterations. In contrast, [Zanelli2021, Thm. 25] considers more general Lyapunov functions, allows for any Lipschitz continuous map $\bar{p}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{p}}$ such that $\bar{u}^{\star}(0)=M_{u,p}\bar{p}(x)$, and considers $k_{\mathrm{max}}=1$. Note that the extension to $k_{\mathrm{max}}>1$ is straight forward because of the q-linear optimizer convergence in Assumption 3. $\square$
IV Decentralized Sequential Quadratic Programming
The crucial requirement on the optimization method to guarantee closed-loop stability via Lemma 1 is local q-linear convergence. Hence we now recall dSQP from [Stomberg2022a] and derive the required convergence property.
By introducing state trajectory copies for neighboring subsystems, OCP (1) can be written as a partially separable Nonlinear Program (NLP) [Summers2012, Bestler2019]
@@ IV Decentralized Sequential Quadratic Programming
subject $\text{subject to}\hskip 1.42262pt\quad g_{i}(z_{i})$ $=0\;|\;\nu_{i}\quad\forall i\in\mathcal{S},$ (5b)
@@
The decision variables $z_{i}\in\mathbb{R}^{n_{i}}$ of subsystem $i$ include the predicted trajectories $\bar{\boldsymbol{x}}_{i}$ and $\bar{\boldsymbol{u}}_{i}$ over the horizon as well as copies of the predicted state trajectories of neighboring subsystems. Example 1 in Appendix B illustrates the reformulation of OCP (2) as a partially separable NLP like (5). The functions $f_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}$, $g_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{g_{i}}}$, and ${h_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{h_{i}}}}$ are composed of the objective functions, equality constraints, and inequality constraints in OCP (1), respectively.
@@
The sparse matrices $E_{i}\in\mathbb{R}^{{n_{c}}\times{n_{i}}}$ couple the subsystems by matching original and copied variables of state trajectories. Thus, the sparsity pattern of the matrices $E_{i}$ arises from the coupling graph $\mathcal{G}$. Specifically, $(i,j)$ or $(j,i)\in\mathcal{E}$ if and only if there exists an index $o\in\{1,\dots,n_{c}\}$ such that $[E_{i}]_{o}\neq 0$ and $[E_{j}]_{o}\neq 0$.
@@
The bi-level dSQP method from [Stomberg2022a] combines an SQP scheme on the outer level with ADMM on the inner level. We index outer iterations by $\cdot^{k}$ and inner iterations by $\cdot^{l}$. Starting from a primal-dual point $p^{k}=(z^{k},\nu^{k},\mu^{k},\lambda^{k})$, the method proceeds as follows. In each SQP iteration, we first construct a quadratic approximation of NLP (5)
@@
subject $\textrm{subject to}\quad g_{i}^{k}+\nabla g_{i}^{k\top}(z_{i}-z_{i}^{k})$ $=0\;|\;\nu_{i}\hskip 9.95845pt\forall i\in\mathcal{S},$ (6b)
@@
where, $f_{i}^{\mathrm{QP},k}\doteq(z_{i}-z_{i}^{k})^{\top}H_{i}^{k}(z_{i}-z_{i}^{k})/2+\nabla f_{i}^{k\top}(z_{i}-z_{i}^{k})$, and where, for all $i\in\mathcal{S}$, $H_{i}^{k}\approx\nabla_{z_{i}z_{i}}^{2}L_{i}(z_{i}^{k},\nu_{i}^{k},\mu_{i}^{k})$ is positive definite on the space spanned by (6b). The symbols $g_{i}^{k}$ and $\nabla g_{i}^{k}$ in QP (6) are shorthands for $g_{i}(z_{i}^{k})$ and $\nabla g_{i}(z_{i}^{k})$, respectively. The same holds for functions $f_{i}$ and $h_{i}$. We denote the unique centralized primal dual-solution to QP (6) as $p^{k,\star}_{\mathrm{QP}}=(z^{k,\star}_{\mathrm{QP}},\nu^{k,\star}_{\mathrm{QP}},\mu^{k,\star}_{\mathrm{QP}},\lambda^{k,\star}_{\mathrm{QP}})$ and the centralized Hessian as $H^{k}\doteq\mathrm{diag}(H_{1}^{k},\dots,H_{S}^{k})$.
@@
subject to $\textrm{subject to }\quad y_{i}-z_{i}$ $=0\;|\;\gamma_{i}\quad\quad\forall i\in\mathcal{S}.$ (7b)
@@
$\mathbb{Z}_{i}^{k}\doteq\left\{y_{i}\in\mathbb{R}^{n_{i}}\left|\begin{aligned} g_{i}^{k}+\nabla g_{i}^{k\top}(y_{i}-z_{i}^{k})&=0\\ h_{i}^{k}+\nabla h_{i}^{k\top}(y_{i}-z_{i}^{k})&\leq 0\end{aligned}\right.\right\},$
@@
$\mathbb{E}\doteq\left\{z\in\mathbb{R}^{n}\left|\sum_{i\in\mathcal{S}}E_{i}z_{i}=c\right.\right\},$
@@
where $L_{\rho,i}^{k}\doteq f_{i}^{\mathrm{QP},k}(y_{i})+\gamma_{i}^{\top}(y_{i}-z_{i})+\rho\|y_{i}-z_{i}\|_{2}^{2}/2$, and where $\rho>0$ is a penalty parameter. The ADMM iterations in the centralized variables read [Boyd2011]
@@
where the notation in (8a)–(8b) indicates that we update the primal and dual iterates with the primal-dual solution obtained in the respective step, and where $\mathbb{Z}^{k}\doteq\mathbb{Z}_{1}^{k}\times\dots\times\mathbb{Z}_{S}^{k}$. ADMM is given in Algorithm 1.
@@
Algorithm 1 ADMM for solving QP (6[Boyd2011]
1:Initialization: $z_{i}^{0},\gamma_{i}^{0}$ for all $i\in\mathcal{S}$, $l_{\text{max}}$
@@
Algorithm 2 dSQP for solving NLP (5[Stomberg2022a]
@@
Remark 2 (Decentralized ADMM)
Steps 3 and 5 of Algorithm 1 can be carried out by each subsystem individually. For the considered OCPs, each constraint in (5d) couples exactly two subsystems to match original and copied state variables. Moreover, $c=0$. Note that every NLP can be transformed to have these so-called 2-assigned constraints by adding further decision variables [Engelmann2020c]. Thus, NLP (5) is a consensus optimization problem and ADMM can be decentralized [Boyd2011, Ch. 7]. Appendix B illustrates the decentralization of the averaging step (8b) on a small-scale example. More details specific to DMPC, including open-source code, are given in [Bestler2019, Stomberg2023, Stomberg2024b]. Following [Nedic2018], we therefore refer to the employed ADMM variant and in consequence to dSQP as decentralized optimization methods.  $\square$
In the following, we combine convergence results from inexact SQP schemes for the outer level and from ADMM for the inner level to pave the road towards distributed RTI schemes with stability guarantees via novel q-linear convergence guarantees for Algorithm 2.
IV-A Outer Convergence
We first consider a basic SQP method, where QP (6) is solved to high accuracy in each SQP iteration [Boggs1995]. Then, we move on to inexact SQP schemes which use approximate solutions of QP (6).

### v3

@@ 3 Centralized Real-Time Iterations
Furthermore, there exists a constant $L_{V,x}>0$ such that $|\sqrt{V(x)}-\sqrt{V(x^{\prime})}|\leq L_{V,x}\|x-x^{\prime}\|$ for all $x,x^{\prime}\in\mathbb{X}_{\bar{V}}$. Moreover, there exists a constant $\hat{r}_{x}>0$ such that $\mathbb{X}_{\bar{V}}\oplus\mathcal{B}(0,\hat{r}_{x})\subseteq\mathbb{X}_{0}$. $\square$
@@
Let $\bar{p}:\mathbb{X}_{0}\rightarrow\mathbb{R}^{n_{p}}$ be the map from the current state $x(t)$ in (2c) to a globally optimal primal-dual variable $p^{\star}$ of OCP (2). $\square$
@@
Assumption 2 (Lipschitz controller [7])
There exists a positive constant $L_{p,x}$ such that $\|\bar{p}(x^{\prime})-\bar{p}(x)\|\leq L_{p,x}\|x^{\prime}-x\|$ for all $x\in\mathbb{X}_{\bar{V}}$ and for all $x^{\prime}\in\mathbb{X}_{\bar{V}}\cup\mathcal{B}(x,\hat{r}_{x})$. Moreover, $\bar{p}(0)=0$. $\square$
Note that $\bar{p}(x^{\prime})$ in Assumption 2 is well defined for all $x^{\prime}\in\mathbb{X}_{\bar{V}}\cup\mathcal{B}(x,\hat{r}_{x})$, because Assumption 1 ensures that OCP (2) with initial state $x^{\prime}$ is feasible. Further note that $\bar{p}(0)=0$ if the origin is an equilibrium of the system dynamics which minimizes $\ell$ and $V_{\mathrm{f}}$, a common design choice in stabilizing NMPC. Put differently, this assumption does not hold for so-called economic NMPC schemes without further modification.
We continue by introducing the considered RTI scheme and the resulting system-optimizer dynmics. Let the superscript $\cdot^{k}$ denote the iteration index of the optimization method which is used to solve OCP (2). At time $t$, the state $x(t)$ is sampled and the optimization method is initialized with the solution from the previous control step, i.e., $p^{0}(t)=p^{k_{\text{max}}}(t-1)$, where $k_{\text{max}}\in\mathbb{N}$ is the number of optimizer iterations per control step. Then, $k_{\text{max}}$ optimizer iterations are applied to OCP (2), a primal-dual iterate $p^{k_{\text{max}}}(t)$ is obtained, and the control input is selected, i.e. $u(t)=M_{u,p}p^{k_{\mathrm{max}}}(t)$. The control input is applied to the system and the primal-dual iterates are stored for the next time step. The optimization method and the system together form the system-optimizer dynamics [7]
$\begin{bmatrix}x(t+1)\\ p^{k_{\mathrm{max}}}(t+1)\end{bmatrix}=\begin{bmatrix}f^{\delta}(x(t),M_{u,p}p^{k_{\mathrm{max}}}(t))\\ \Phi(x(t),p^{k_{\text{max}}}(t))\end{bmatrix}$, (4)
where $\Phi:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{p}}\rightarrow\mathbb{R}^{n_{p}}$ maps the OCP initial state $x(t)$ and approximate solution $p^{k_{\text{max}}}(t)$ to the approximate solution $p^{k_{\text{max}}}(t+1)$ at the next time step. We tailor [7, Ass. 10] by assuming optimizer convergence to the current OCP solution $\bar{p}(x(t))$ as follows.
Assumption 3 (Q-linear optimizer convergence)
There exist positive constants $\hat{r}_{p}>0$ and $a_{p}<1$ such that, for all $x(t)\in\mathbb{X}_{\bar{V}}$ and all $p^{0}(t)\in\mathcal{B}(\bar{p}(x(t)),\hat{r}_{p})$, the sequence $\{p^{k}(t)\}$ of optimizer iterates satisfies
$\left\|p^{k+1}(t)-\bar{p}(x(t))\right\|\leq a_{p}\left\|p^{k}(t)-\bar{p}(x(t))\right\|\;\forall k\in\mathbb{N}_{0}.$
@@
Assumption 4 (Lipschitz system dynamics [7])
The centralized dynamics satisfy $f^{\mathrm{c}}(0,0)=0$ and there exist positive finite constants $r_{p}^{\prime}$, $\varrho$, $L_{f,x}^{\mathrm{c}}$, and $L_{f,u}^{\mathrm{c}}$ such that
@@
for all $x^{\prime},x\in\mathbb{X}_{\bar{V}}\oplus\mathcal{B}(0,\varrho)$ and for all $u^{\prime}=M_{u,p}p^{\prime}$, $u=M_{u,p}p$ with $p^{\prime},p\in\mathcal{B}(\bar{p}(x),r_{p}^{\prime})$. $\square$
Lemma 1 (Centralized RTI stability)
Suppose that As-sumptions 14 hold and consider the sufficient sampling interval $\bar{\delta}$ and optimizer initialization radius $\tilde{r}_{p}$ defined in (31) in Appendix 8. If $\delta\leq\bar{\delta}$, then the origin is a locally exponentially stable equilibrium with region of attraction
[…]
for the closed-loop system-optimizer dynamics (4). $\square$
Proof 3.1.
The proof proceeds similarly to the analysis in [7]. Since our q-linear optimizer convergence Assumption 3 differs from [7, Ass. 10], we replace [7, Lem. 11] by Lemma 22 from the Appendix. Specifically, the optimizer contraction inequality (30) derived in Lemma 22 is equivalent to [7, Ineq. 26] in the proof of [7, Prop. 16]. The remaining analysis in [7] stays applicable such that local exponential stability follows from [7, Thm. 25].
Remark 1 (Relation to [7]).
Lemma 1 is a specialized variant of [7, Thm. 25] in four aspects: First, we only consider quadratic bounds on the Lyapunov function. Second, we define the map $\bar{p}$ from the current state to the primal-dual variables, because of the dSQP convergence presented in the next section. Third, we allow for $k_{\mathrm{max}}\geq 1$ optimizer iterations. In contrast, [7, Thm. 25] considers more general Lyapunov functions, allows for any Lipschitz continuous map $\bar{p}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{p}}$ such that $\bar{u}^{\star}(0)=M_{u,p}\bar{p}(x)$, and considers $k_{\mathrm{max}}=1$. Note that the extension to $k_{\mathrm{max}}>1$ is straight forward because of the q-linear optimizer convergence in Assumption 3. Moreover, the rationale of this section is to first select a candidate sampling interval $\delta$ such that Assumptions 14 hold and to then infer stability if $\delta\leq\bar{\delta}$. The approach of [7, Thm. 25] instead lets an assumption similar to Assumption 1 hold for a range of sampling intervals and then guarantees stability for all sampling intervals below $\bar{\delta}$. $\square$
4 Decentralized Sequential Quadratic Programming
The crucial requirement on the optimization method to guarantee closed-loop stability via Lemma 1 is local q-linear convergence. Hence we now recall dSQP from [28] and derive the required convergence property.
By introducing state trajectory copies for neighboring subsystems, OCP (1) can be written as a partially separable Nonlinear Program (NLP) [18, 16]
@@ 4 Decentralized Sequential Quadratic Programming
$\text{subject to}\hskip 1.42262pt\quad g_{i}(z_{i})$ $=0\;|\;\nu_{i}\quad\forall i\in\mathcal{S},$ (5b)
@@
The decision variables $z_{i}\hskip-2.84544pt\in\hskip-2.84544pt\mathbb{R}^{n_{i}}$ of subsystem $i$ include the predicted trajectories $\bar{\boldsymbol{x}}_{i}$ and $\bar{\boldsymbol{u}}_{i}$ over the horizon as well as copies of the predicted state trajectories of neighboring subsystems. Example 9.1 in Appendix 9 illustrates the reformulation of OCP (1) as a partially separable NLP like (5). The functions $f_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}$, $g_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{g_{i}}}$, and $h_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{h_{i}}}$} are composed of the objective functions, equality constraints, and inequality constraints in OCP (1), respectively.
@@
The sparse matrices $E_{i}\in\mathbb{R}^{{n_{c}}\times{n_{i}}}$ couple the subsystems by matching original and copied variables of state trajectories. Thus, the sparsity pattern of the matrices $E_{i}$ arises from the coupling graph $\mathcal{G}$. Specifically, $(i,j)$ or $(j,i)\in\mathcal{E}$ if and only if there exists an index $o\in\{1,\dots,n_{c}\}$ such that $E_{i}]_{o}\neq 0$ and $E_{j}]_{o}\neq 0$.
@@
The bi-level dSQP method from [28] combines an SQP scheme on the outer level with ADMM on the inner level. We index outer iterations by $\cdot^{k}$ and inner iterations by $\cdot^{l}$. Starting from a primal-dual point $p^{k}=(z^{k},\nu^{k},\mu^{k},\lambda^{k})$, the method proceeds as follows. In each SQP iteration, we first construct a quadratic approximation of NLP (5)
@@
$\textrm{subject to}\quad g_{i}^{k}+\nabla g_{i}^{k\top}(z_{i}-z_{i}^{k})$ $=0\;|\;\nu_{i}\hskip 9.95845pt\forall i\in\mathcal{S},$ (6b)
@@
[…]
@@
$\textrm{subject to }\quad y_{i}-z_{i}$ $=0\;|\;\gamma_{i}\quad\quad\forall i\in\mathcal{S}.$ (7b)
@@
[…]
@@
[…]
@@
where $L_{\rho,i}^{k}\doteq f_{i}^{\mathrm{QP},k}(y_{i})+\gamma_{i}^{\top}(y_{i}-z_{i})+\rho\|y_{i}-z_{i}\|_{2}^{2}/2$ and where $\rho>0$ is a penalty parameter. The ADMM iterations in the centralized variables read [34]
@@
where the notation in (8a)–(8b) indicates that we update the primal and dual iterates with the primal-dual solution obtained in the respective step and where $\mathbb{Z}^{k}\doteq\mathbb{Z}_{1}^{k}\times\dots\times\mathbb{Z}_{S}^{k}$. ADMM is given in Algorithm 1.
@@
Algorithm 1 ADMM for solving QP (6[34]
1:Initialization: $z_{i}^{0},\gamma_{i}^{0}$, $l_{\text{max}}$ for all $i\in\mathcal{S}$,
@@
Algorithm 2 dSQP for solving NLP (5[28]
@@
Remark 2 (Decentralized ADMM).
Steps 3 and 5 of Algorithm 1 can be carried out by each subsystem individually. For the considered OCPs, each constraint in (5d) couples exactly two subsystems to match original and copied state variables. Moreover, $c=0$. Note that every NLP can be transformed to have these so-called 2-assigned constraints by adding further decision variables [35]. Thus, NLP (5) is a consensus optimization problem and ADMM can be decentralized [34, Ch. 7]. Appendix 9 illustrates the decentralization of (8b) on a small-scale example. Following [22], we therefore refer to the employed ADMM variant and in consequence to dSQP as decentralized optimization methods. $\square$
Remark 3 (Communication requirements).
The only step of the proposed dRTI scheme which requires communication is Step 4 of ADMM. As discussed in Remark 2, this step can be implemented as an efficient averaging step which only requires neighbor-to-neighbor communication. A common procedure considered in DMPC is to implement the averaging step via two communication rounds: First, copied state trajectories are sent to in-neighbors. Then, averaged state trajectories are sent to out-neighbors, cf. [16, 29, 30]. $\square$
In the following, we combine convergence results from inexact SQP schemes for the outer level and from ADMM for the inner level to pave the road towards decentralized RTI schemes with stability guarantees via novel q-linear convergence guarantees for Algorithm 2.
4.1 Outer Convergence
We first consider a basic SQP method, where QP (6) is solved to high accuracy in each SQP iteration [36]. Then, we move on to inexact SQP schemes which use approximate solutions of QP (6).